OpenStudy (anonymous):

The function f(x) = 245(4)x represents the growth of a fruit fly population every year in the orange grove. Stacey wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Stacey's purposes?

OpenStudy (anonymous):

f(x) = 45(4)x f(x) = 45(412)the x over 12 power f(x) = 2752(4)x f(x) = 245(4 to the one twelfth power)12x are my choices

OpenStudy (anonymous):

U have an idea

OpenStudy (anonymous):

no im not sure where to begin

OpenStudy (anonymous):

@Agl202 thanks so much im still confused on which awnser choice it would be:(

OpenStudy (anonymous):

I'm pretty sure you can get rid of a and c

OpenStudy (agl202):

Im pretty sure the correct answer choice would be D.

OpenStudy (anonymous):

ok! thanks soo much

OpenStudy (agl202):

Ur welcome!